Find the phase speed v for the wave whose snapshot and history graphs are given below. Show your calculations for the wave crest. [ans. -1.5 m/s]

Please show work :)
5. [5pt Find the phase speed v for the wave whose snapshot and history graphs are given below. Show your calculations for the wave crest. [ans. -1.5 m/s] 3-1 211 t = 8 0 1 2 3 5 6 78 9 10 11 12 X (m) 2x7 m -3 0 1 23 5 678 9 10 11 12 t (s) Question to be Graded Manually.
Answer #1

From First graph:

* Length of pulse = wavelength = 10 - 4 = 6 m

From Second graph:

* Time period of pulse = 10 - 6 = 4 s

phase speed = wavelength/ time period = 6 / 4 = 1.5 m/s